OpenStudy (anonymous):

I WILL GIVE MEDAL Solve and Check c/3 = 6 1/7

OpenStudy (anonymous):

@meeya Please help ill give u a medal

OpenStudy (anonymous):

Please Help me

OpenStudy (anonymous):

do you know wat c is

OpenStudy (anonymous):

Its a variable isnt it? and no thats what we are trying to figure out

OpenStudy (anonymous):

im learning that to so i dont really know

OpenStudy (anonymous):

oh thanks anyway :D

OpenStudy (anonymous):

this is wat im doing 5x-3=17

OpenStudy (anonymous):

oh I can help you i think

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I think i figured out the answer to it, so here goes

OpenStudy (anonymous):

So do this first, 17+3

OpenStudy (anonymous):

So what is the answer to that?

OpenStudy (anonymous):

-8b+10+10b-5-2x

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

it anther question

OpenStudy (anonymous):

oh, lets do the first one first.

OpenStudy (anonymous):

17+3=20

OpenStudy (anonymous):

ok, so ,

OpenStudy (anonymous):

if thats 20, then 20-3 would be 17,

OpenStudy (anonymous):

so if 20-3 equals 17, and 17 is the sum of the equasion, what is x, in 5x?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

5x would be 20, so x is 4, 5x4=20-3=17

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

by what?

OpenStudy (anonymous):

that

OpenStudy (anonymous):

what?